# Volatile-rich Sub-Neptunes as Hydrothermal Worlds: The Case of K2-18 b

### v1

@@ 3.1 Can a Supercritical Water Ocean Explain K2-18 b’s Atmospheric Chemistry?
Our modeling approach examines both the aqueous and gas phases to evaluate the impact of phase transitions on the ratios of carbon species in K2-18 b’s hypothetical supercritical water ocean and in its atmosphere. However, we focus primarily on the gaseous ratios results, since they correspond most closely to the observed abundances in K2-18 b’s upper atmosphere. For the CH4/CO2 ratio (Figure 3), our results show that higher ocean temperatures favor the production of CO2, while higher pressures and higher H2 molalities promote the formation of CH4. We find that the predicted gaseous CH4/CO2 ratios indicate a lower temperature limit of T $\approx$ 713 K to reproduce the observed CH4/CO2 ratio, across the full range of H2 molalities explored. Sub-critical temperatures would only be consistent with the measured CH4/CO2 ratio if the H2 abundance were very low, requiring K2-18 b to have an unrealistically high water/rock ratio. For the CO/CO2 ratio (Figure 4), we see that higher temperatures and lower pressures favor the formation of CO. We find that our gaseous CO/CO2 ratios support temperatures less than $\sim$ 1070 K to avoid exceeding the observational detection limit for CO (Madhusudhan et al., 2023b).
@@
Combining these constraints then provides a self-consistent ocean temperature range of T $\approx$ 713 K – 1070 K, which exceeds the critical point of water. These findings suggest that the presence of a supercritical water ocean can explain the observed carbon ratios in K2-18 b’s atmosphere. The presence of such an ocean, combined with water vapor cold-trapped in its observable photosphere, would categorize K2-18 b as a “Stratified Mini-Neptune” as defined by Benneke et al. (2024). Note that unlike “mini-Neptune” gas-phase models that assume deep H2-dominated atmospheres for K2-18 b (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Leconte et al., 2024; Wogan et al., 2024; Yang & Hu, 2024), our supercritical ocean chemistry model can explain the low observed abundance of CO on K2-18 b.
@@
Both the lower and upper temperature limits are conservative estimates. Higher pressures than those we explored are possible, and if equilibrium is attained at higher pressures, this may shift the lower limit on temperature upwards. Meanwhile, CO-CO2 re-equilibrates faster than CH4-CO2 since these reactions require more bond-breaking and bond-making steps (Wei & Iglesia, 2004), and thus CO may re-equilibrate in the warm atmosphere above the ocean.
@@
Figure 4: Aqueous and gaseous CO/CO2 ratios, the symbols and the colors are labeled the same ways as in Figure 3. The region of error surrounding the curves is due to the one-offset error for the CH4/CO2 ratio.
@@
We note that the gaseous results show only subtle variations compared to the aqueous phase, for both the CH4/CO2 and CO/CO2 ratios, which align with our expectations, given the relatively similar solubilities of CH4, CO2, and CO in supercritical water. CO2 is slightly more soluble in liquid water than CH4 and CO are, which may cause some fractionation during water cloud condensation. However, this effect diminishes with increasing temperature. As the temperature rises, their respective Henry’s Law constants become more similar (see Figure F.1). This reduces their differences in solubilities and minimizes fractionation. As a result, the ratios in the gaseous phase closely mirror those of the aqueous phase, with only minor variations driven by the slightly higher solubility of CO2.
@@
Although our gaseous calculations represent carbon ratios at the cloud interface layer, we can directly compare these results with observed abundances from the region probed by JWST. We assume the gaseous ratios remain largely unchanged for two main reasons. First, previous photochemical models suggest that these carbon species are quenched at pressures between 10-100 bar (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Yang & Hu, 2024), similar to the pressure levels at the base of the cloud layer where our calculations are performed. At pressures lower than the quench point, the mixing ratios of these species remain constant as they are transported upward through the atmosphere until photochemistry becomes important. Since JWST primarily probes the atmosphere at pressures between 0.1-2 mbar (Rustamkulov et al., 2023), and photochemistry dominates for these carbon species higher in the atmosphere at pressures of 0.1 mbar or less (Hu et al., 2021; Tsai et al., 2021; Yu et al., 2021; Wogan et al., 2024), our calculations are well-suited for comparison with JWST data. Secondly, the relative insolubility of these carbon species in liquid water clouds suggests that their gaseous ratios should undergo minimal change as they move through the water cloud. As a result, our calculated gaseous ratios at the interface with the cloud layer can be directly compared with JWST observations.
@@ 3.2 Implications for Habitability
A key question surrounding K2-18 b is whether it can host a habitable environment. Atmospheric observations suggest the ocean temperature should be at least 713 K to explain the observed carbon ratios, assuming thermochemical equilibrium controls dissolved gas abundances in a supercritical water ocean. Currently, the theoretical upper-temperature limit for microbial growth is considered to be $\sim$ 150C (423.15 K) (Schulze-Makuch et al., 2017), indicating that a supercritical water ocean cannot support life as we know it. However, life existing solely within the clouds is a possibility and has been explored for sub-Neptunes by Seager et al. (2021). Although a supercritical water ocean may not permit habitability, studying its dynamics offers a unique opportunity to understand exotic environments on exoplanets like K2-18 b.
A global supercritical water ocean could foster a hydrothermal environment characterized by hot aqueous fluids. Supercritical fluids have been observed in various hydrothermal vent systems on Earth (Von Damm et al., 2003; Koschinsky et al., 2008), however, we are not implying the presence of individual hydrothermal vents on a rocky seafloor on K2-18 b. Instead, K2-18 b may be a planetary-scale hydrothermal system unlike anything known on Earth. Beyond our planet, hydrothermal activity has only been implicated by data from Saturn’s moon Enceladus (Hsu et al., 2015; Waite et al., 2017), and hypothesized on Jupiter’s moon Europa (Melwani Daswani et al., 2021; Běhounková et al., 2021), with no evidence found on an exoplanet yet. While our study does not confirm the existence of hydrothermal processes on K2-18 b, it motivates further investigation into this possibility. The methodology developed here can serve as a first step to investigating hydrothermal planets.
@@ 3.3 Future Work
Our result is consistent with previous climate models of sub-Neptunes, which suggest that water at the lower boundary would likely exist in a supercritical state rather than a liquid state (Innes et al., 2023; Pierrehumbert, 2023; Leconte et al., 2024). Having outlined constraints for a hypothetical supercritical water ocean on K2-18 b, we now turn to the next steps toward an improved understanding of its interior using atmospheric observations.
The most well-constrained abundances from current JWST observations are those of CH4 and CO2, which yield a CH4/CO2 ratio on the order of unity (Madhusudhan et al., 2023b). As these measurements are refined with new data, shifts in this ratio within the error range could occur but may still support the possibility of a supercritical water ocean. Current atmospheric models consistent with the CH4 and CO2 observations include a thin-atmosphere liquid water world with biological CH4 flux, a mini-Neptune lacking a surface influence, a mini-Neptune with a silicate magma ocean, and the supercritical hydrothermal water world modeled here. To distinguish between these possibilities, the abundances of CO and NH3 then become critical factors, as each model predicts different levels of these species (Shorttle et al., 2024; Wogan et al., 2024). Our supercritical water ocean model spans a wide range of CO values, but NH3 was not considered in this study due to the additional complexities involved in modeling its abundances. Factors such as ocean pH, total nitrogen inventory, and NH3 dissolution in water clouds introduce greater uncertainties that make NH3 more challenging to model compared to the carbon species.
While current models have brought us closer to understanding K2-18 b’s atmospheric structure, more precise constraints on NH3 and CO abundances will be instrumental for additional testing and refining of our models. Ongoing and future JWST observations (JWST C1 GO 2372; PI: Renyu Hu, JWST C1 GO 2722; PI: Nikku Madhusudhan) will be crucial in providing these constraints, helping to distinguish between the candidate atmosphere-interior structures.
@@ 4 Conclusion
Volatile-rich sub-Neptunes have sparked intrigue, with their observed masses and radii hinting at a range of potential interior structures, each carrying distinct implications for the planet’s interior processes. Recently, transmission spectroscopy of K2-18 b revealed the first detection of CH4 and CO2 in the atmosphere of a sub-Neptune exoplanet, along with a nondetection of CO. These data allow us to evaluate the consistency of various structural models against the observed atmospheric composition. In this work, we explored the possibility of K2-18 b being a global hydrothermal planet and found that:
  • A supercritical water ocean with temperatures ranging from approximately 713 K to 1070 K and pressures between 1 and 10 kbar can explain the CH4/CO2 ratio observed by JWST and the nondetection of CO.
@@
  • The presence of a supercritical water ocean implies that K2-18 b may be a global hydrothermal environment.
Overall, this work corroborates the possibility of K2-18 b being a hydrothermal world and explores Earth-based geochemical modeling as a valuable tool for characterizing exoplanet interiors. Future observations that better constrain the abundances of NH3 and CO will be critical for advancing our understanding of K2-18 b and other temperate sub-Neptunes.
@@ 5 Acknowledgements
X. Yu, C. N. Luu, J. I. Moses, and J. Krissansen-Totton are supported by the NASA Habitable Worlds Program Grant 80NSSC24K0075. C. R. Glein and N. Truong acknowledge support from the Heising-Simons Foundation (grant 2023-4657). X. Yu is supported by the NASA Planetary Science Early Career Award 80NSSC23K1108 and by the Heising-Simons Foundation grant 2023-3936. S.-M.T. is supported by NASA through Exobiology Grant No. 80NSSC20K1437 and Interdisciplinary Consortia for Astrobiology Research (ICAR) Grant Nos. 80NSSC23K1399, 80NSSC21K0905, and 80NSSC23K1398. X. Zhang is supported by the National Science Foundation grant AST2307463 and the NASA Exoplanet Research grant 80NSSC22K0236. A. Aguichine is supported by NASA’s Interdisciplinary Consortia for Astrobiology Research (ICAR) Grant No. 80NSSC21K0597. This work also benefited from the Exoplanet Summer Programs in the Other Worlds Laboratory (OWL) at the University of California, Santa Cruz, a program funded by the Heising-Simons Foundation. We also acknowledge Dr. Ligia Fonseca Coelho for insightful discussions on the temperature range for microbial growth.
@@ Appendix A Observational data selection
In this paper, we compare our model predictions against abundances computed using the retrieved one-offset mixing ratio values listed in Table 2 of Madhusudhan et al. (2023b). The offset solutions are preferred as numerous studies have suggested the existence of an offset between the NRS1 and NRS2 detectors of the Near-Infrared Spectrograph (NIRSpec) instrument used to observe K2-18 b (May et al., 2023; Moran et al., 2023). We propagate the uncertainty of the observed ratios using the largest mixing ratio error given for each species.
@@ Appendix B Conversion Between Mass Fraction and Hydrogen Molality
To convert the mass fraction percentages into hydrogen molalities in Section 2.2, we refer to solar metallicity values for the H-He mixture consisting of $X_{0}=73.90$% H and $Y_{0}=24.69$% He (Lodders et al., 2009). Here, $X_{0}$ and $Y_{0}$ refer to protosolar mass fractions. If purely H-He are accreted from the nebula, while excluding all silicate dust and volatile ices, the composition of such mixture is 74.96% H and 25.04% He by weight. For a planet with a bulk composition determined by the H-He, H2O and rocky core mass fractions $f_{\mathrm{H-He}}$, $f_{\mathrm{H_{2}O}}$ and $f_{\mathrm{Rock}}=1-f_{\mathrm{H-He}}-f_{\mathrm{H_{2}O}}$, respectively, the molality can be computed as follows:
@@
where $\mu_{\mathrm{H_{2}}}$ is the molar mass of H2 in kg/mol.

### v2

@@ 3.1 Can a Supercritical Water Ocean Explain K2-18 b’s Atmospheric Chemistry?
Our modeling approach examines both the aqueous and gas phases to evaluate the impact of phase transitions on the ratios of carbon species in K2-18 b’s hypothetical supercritical water ocean and in its atmosphere. However, we focus primarily on the results for the gaseous ratios since they correspond most closely to the observed abundances in K2-18 b’s upper atmosphere. For the CH4/CO2 ratio (Figure 3), our results show that higher ocean temperatures favor the production of CO2, while higher pressures and lower water mole fractions (more H2) promote the formation of CH4. We find that the predicted gaseous CH4/CO2 ratios indicate a lower temperature limit of T $\approx$ 710 K to reproduce the observed CH4/CO2 ratio, across the full range of H2 molalities explored. Sub-critical temperatures would only be consistent with the measured CH4/CO2 ratio if the H2 abundance were very low, requiring K2-18 b to have an unrealistically high water/rock mass ratio. For the CO/CO2 ratio (Figure 4), it can be seen that higher temperatures and lower pressures favor the formation of CO. We find that our gaseous CO/CO2 ratios support temperatures less than $\sim$ 1070 K to avoid exceeding the observational detection limit for ${}_{\text{CO}}$) $<-3.50$ at a 95% upper confidence limit, (Madhusudhan et al., 2023b)).
@@
Combining these constraints then provides a self-consistent ocean temperature range of T $\approx$ 710 K – 1070 K, which exceeds the critical temperatures of pure water as well as relevant H2-H2O mixtures (Seward & Franck, 1981). These findings suggest that the presence of a supercritical water ocean can explain the observed carbon ratios in K2-18 b’s atmosphere. Between pressures of 1 - 5 kbar, H2 is completely miscible with water according to Gupta et al. (2024), leading to a homogeneous supercritical H2-H2O ocean that would categorize K2-18 b as a “Stratified Mini-Neptune” as defined by Benneke et al. (2024). On the other hand, at pressures between 5 - 10 kbar, H2 and H2O may be immiscible at some temperatures, leading to a phase-separated water-rich layer beneath a supercritical H2-rich layer (Gupta et al., 2024). We discuss more intricate scenarios and their implications further in Appendix G.3. This structure would categorize K2-18 b as a mini-Neptune similar to Uranus and Neptune interior models (Scheibe et al., 2019; Gupta et al., 2024; Cano Amoros et al., 2024). Note that unlike “mini-Neptune” gas-phase models that assume deep H2-dominated atmospheres for K2-18 b (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Leconte et al., 2024; Yang & Hu, 2024), our supercritical ocean chemistry model can explain the low observed abundance of CO on K2-18 b.
@@
Both the lower and upper temperature limits are conservative estimates. Higher pressures than those we explored are possible, and if equilibrium is attained at higher pressures, this may shift the lower limit on temperature upwards. In addition, the CO-CO2 couple re-equilibrates faster than CH4-CO2 since the latter reaction requires more bond-breaking and bond-making steps (Wei & Iglesia, 2004), and thus CO may partially re-equilibrate in the warm atmosphere above the ocean.
@@
Figure 4: (a-c) Aqueous and (d-f) gaseous CO/CO2 ratios. The symbols and the colors are labeled the same ways as in Figure 3. The region of uncertainty surrounding the curves is due to the one-offset error for the CH4/CO2 ratio.
@@
We note that the gaseous results show only subtle differences compared to the aqueous phase, for both the CH4/CO2 and CO/CO2 ratios, which align with our expectations, given the relatively similar solubilities of CH4, CO2, and CO in supercritical water. CO2 is slightly more soluble in liquid water than CH4 and CO are, which may cause some fractionation during water cloud condensation. However, this effect diminishes with increasing temperature. As the temperature rises, their respective Henry’s Law constants become more similar (see Figure F.1). This reduces their differences in solubility and minimizes fractionation. As a result, the ratios in the gaseous phase closely mirror those of the aqueous phase, with only minor variations driven by the slightly higher solubility of CO2.
@@
Although our gaseous calculations represent carbon ratios at the cloud interface layer, we can directly compare these results with observed abundances from the region probed by JWST. We assume the gaseous ratios remain largely unchanged for two main reasons. First, previous photochemical models suggest that these carbon species are quenched at pressures between 10-100 bar (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Yang & Hu, 2024), similar to the pressure levels at the base of the cloud layer where our calculations are performed. At pressures lower than the quench point, the mixing ratios of these species remain constant as they are transported upward through the atmosphere until photochemistry becomes important. Since JWST primarily probes the atmosphere at pressures between 0.1-2 mbar (Rustamkulov et al., 2023) and multiple photochemical models demonstrate that photochemistry may have minimal influence on the atmospheric abundances within this region (Tsai et al., 2021; Yu et al., 2021; Tsai et al., 2024; Huang et al., 2024; Yang & Hu, 2024), our calculations are well-suited for comparison with JWST data. Secondly, the relative insolubility of these carbon species in liquid water clouds suggests that their gaseous ratios should undergo minimal change as they move through the water cloud. As a result, our calculated gaseous ratios at the interface with the cloud layer can be directly compared with JWST observations.
@@ 3.2 Implications for Habitability
A key question surrounding K2-18 b is whether it can host a habitable environment. Atmospheric observations suggest the ocean temperature should be at least 710 K to explain the observed carbon ratios, assuming thermochemical equilibrium controls dissolved gas abundances in a supercritical water ocean. Currently, the theoretical upper-temperature limit for microbial growth is considered to be $\sim$ 150C (423.15 K) (Schulze-Makuch et al., 2017), indicating that a supercritical water ocean cannot support life as we know it. However, life existing solely within the clouds is a possibility and has been explored for sub-Neptunes by Seager et al. (2021). Although a supercritical water ocean may not permit habitability, studying its chemistry offers a unique opportunity to understand exotic environments on exoplanets like K2-18 b.
A global supercritical water ocean could foster a hydrothermal environment characterized by hot aqueous fluids. Supercritical fluids have been observed in hydrothermal vent systems on Earth (Von Damm et al., 2003; Koschinsky et al., 2008), however, we are not implying the presence of individual hydrothermal vents on a rocky seafloor on K2-18 b. Instead, K2-18 b may be a planetary-scale hydrothermal system unlike anything known on Earth. Beyond our planet, similar hydrothermal activity has only been implicated by data from Saturn’s moon Enceladus (Hsu et al., 2015; Waite et al., 2017), and hypothesized on Jupiter’s moon Europa (Melwani Daswani et al., 2021; Běhounková et al., 2021), with no evidence found on an exoplanet to date. While our study does not confirm the existence of hydrothermal processes on K2-18 b, it motivates further investigation into this possibility. The methodology developed here can serve as a first step to investigating hydrothermal worlds.
@@ 3.3 Future Work
Our finding is consistent with previous climate models of sub-Neptunes, which suggest that water at the lower boundary would likely exist in a supercritical state rather than a liquid state (Innes et al., 2023; Pierrehumbert, 2023; Leconte et al., 2024). Having outlined constraints for a hypothetical supercritical water ocean on K2-18 b, we now turn to the next steps toward an improved understanding of its interior using atmospheric observations.
The most well-constrained abundances from current JWST observations are those of CH4 and CO2, which yield a CH4/CO2 ratio on the order of unity (Madhusudhan et al., 2023b). As these measurements are refined with new data, shifts in this ratio within the error range could occur but may still support the possibility of a supercritical water ocean. Current atmospheric models consistent with the CH4 and CO2 observations include a thin-atmosphere liquid water world with biological CH4 flux, a mini-Neptune lacking a surface influence, and the supercritical hydrothermal water world modeled here. To distinguish between these possibilities, the abundances of CO and NH3 then become critical factors, as each model predicts different levels of these species (Shorttle et al., 2024; Wogan et al., 2024; Yang & Hu, 2024). Our supercritical water ocean model spans a wide range of CO/CO2 ratios, but NH3 was not considered in this study due to the additional complexities involved in modeling its geochemistry. Factors such as ocean pH, total nitrogen inventory, and NH3 dissolution in water clouds introduce greater uncertainties that make NH3 more challenging to model compared to the carbon species. Continued study of the phase behavior of H2-H2O mixtures under conditions relevant to sub-Neptunes and the development of geochemically practical parameterizations would also be useful next steps.
While current models have brought us closer to understanding K2-18 b’s atmospheric structure, more precise constraints on CO and NH3 abundances will be instrumental for additional testing and refining of our models. Ongoing and future JWST observations (JWST C1 GO 2372; PI: Renyu Hu, JWST C1 GO 2722; PI: Nikku Madhusudhan) will be crucial in providing these constraints, helping to distinguish between the candidate atmosphere-interior structures.
@@ 4 Conclusion
Volatile-rich sub-Neptunes have sparked intrigue, with their observed masses and radii hinting at a range of potential interior structures, each carrying distinct implications for the planet’s inner workings and the origin of these planets. Recently, transmission spectroscopy of K2-18 b revealed the first detection of CH4 and CO2 in the atmosphere of a sub-Neptune exoplanet, along with a nondetection of CO. These data allow us to evaluate the consistency of various structural models against the observed atmospheric composition. In this work, we explored the possibility of K2-18 b being a global hydrothermal world and found that:
  • A supercritical water ocean with temperatures ranging from approximately 710 K to 1070 K and pressures between 1 and 10 kbar can explain the CH4/CO2 ratio observed by JWST and the nondetection of CO.
@@
  • The presence of a supercritical water ocean implies that K2-18 b may host a global hydrothermal environment.
Overall, this work corroborates the possibility of K2-18 b being a hydrothermal world and explores how Earth-based geochemical modeling can be a valuable tool for characterizing exoplanet interiors. Future observations that better constrain the abundances of CO and NH3 will be critical for advancing our understanding of K2-18 b and other temperate sub-Neptunes.
@@ 5 Acknowledgements
X. Yu, C. N. Luu, J. I. Moses, and J. Krissansen-Totton are supported by the NASA Habitable Worlds Program Grant 80NSSC24K0075. C. R. Glein and N. Truong acknowledge support from the Heising-Simons Foundation (grant 2023-4657). X. Yu is supported by the NASA Planetary Science Early Career Award 80NSSC23K1108 and by the Heising-Simons Foundation grant 2023-3936. S.-M.T. is supported by NASA through Exobiology Grant No. 80NSSC20K1437 and Interdisciplinary Consortia for Astrobiology Research (ICAR) Grant Nos. 80NSSC23K1399, 80NSSC21K0905, and 80NSSC23K1398. X. Zhang is supported by the National Science Foundation grant AST2307463 and the NASA Exoplanet Research grant 80NSSC22K0236. A. Aguichine is supported by NASA’s Interdisciplinary Consortia for Astrobiology Research (ICAR) Grant No. 80NSSC21K0597. H.I. is supported by the European Union (ERC, DIVERSE, 101087755). This work also benefited from the Exoplanet Summer Programs in the Other Worlds Laboratory (OWL) at the University of California, Santa Cruz, a program funded by the Heising-Simons Foundation. We also acknowledge Dr. Ligia Fonseca Coelho for insightful discussions on the temperature range for microbial growth.
@@ Appendix A Observational data selection
In this paper, we compare our model predictions against abundances computed using the retrieved one-offset mixing ratio values listed in Table 2 of Madhusudhan et al. (2023b). The offset solutions are preferred as numerous studies have suggested the existence of an offset between the NRS1 and NRS2 detectors of the Near-Infrared Spectrograph (NIRSpec) instrument used to observe K2-18 b (May et al., 2023; Moran et al., 2023). We propagate the uncertainty of the observed ratios using the largest mixing ratio error given for each species. Given the uncertainties in the atmospheric retrieval, it is desirable to adopt the most conservative posture.
@@ Appendix B Conversion Between Mass Fraction and Hydrogen Molality
To convert the mass fraction percentages into hydrogen molalities in Section 2.2, we refer to protosolar metallicity values for the H-He mixture consisting of $X_{0}=70.04$% H and $Y_{0}=27.80$% He (Truong et al., 2024). Here, $X_{0}$ and $Y_{0}$ refer to protosolar mass fractions. If purely H-He are accreted from the protostellar nebula, while excluding all silicate dust and volatile ices, the composition of the gas is 71.58% H and 28.41% He by weight. For a planet with a bulk composition determined by the H-He, H2O and rocky core mass fractions $f_{\mathrm{H-He}}$, $f_{\mathrm{H_{2}O}}$ and $f_{\mathrm{Rock}}=1-f_{\mathrm{H-He}}-f_{\mathrm{H_{2}O}}$, respectively, the molality can be computed as follows:
@@
where $\mu_{\mathrm{H_{2}}}$ stands for the molar mass of H2 in $\cdot$(mol-1).